### v3

# Semantic Equitable Clustering: A Simple and Effective Strategy for Clustering Vision Tokens

Abstract
The Vision Transformer (ViT) has gained prominence for its superior relational modeling prowess. However, its global attention mechanism’s quadratic complexity poses substantial computational burdens. A common remedy spatially groups tokens for self-attention, reducing computational requirements. Nonetheless, this strategy neglects semantic information in tokens, possibly scattering semantically-linked tokens across distinct groups, thus compromising the efficacy of self-attention intended for modeling inter-token dependencies. Motivated by these insights, we introduce a fast and balanced clustering method, named Semantic Equitable Clustering (SEC). SEC clusters tokens based on their global semantic relevance in an efficient, straightforward manner. In contrast to traditional clustering methods requiring multiple iterations, our method achieves token clustering in a single pass. Additionally, SEC regulates the number of tokens per cluster, ensuring a balanced distribution for effective parallel processing on current computational platforms without necessitating further optimization. Capitalizing on SEC, we propose a versatile vision backbone, SECViT. Comprehensive experiments in image classification, object detection, instance segmentation, and semantic segmentation validate the effectiveness of SECViT. Moreover, SEC can be conveniently and swiftly applied to multimodal large language models (MLLM), such as LLaVA, to serve as a vision language connector, effectively accelerating the model’s efficiency while maintaining unchanged or better performance.
1 Introduction
Since its inception, the Vision Transformer (ViT)[11] has drawn considerable interest from the research community due to its robust modeling prowess. However, the quadratic complexity of Self-Attention leads to significant computational overhead, thus constraining the practicality of ViT. A variety of strategies have been devised to alleviate this computational load, the most prevalent of which involves token grouping, thereby constraining the attention span of each token[38, 10, 47, 44].
Specifically, the Swin-Transformer [38] partitions tokens into multiple small windows, restricting token attention within each window. The CSWin-Transformer [10] adopts a cross-shaped grouping, endowing each token with a global receptive field. MaxViT [44] amalgamates window and grid attention, facilitating intra-window tokens to attend to their counterparts in other windows. However, these methods, solely reliant on spatial positioning, neglect token semantics, potentially restricting the self-attention’s capacity to model semantic dependencies. To mitigate this, DGT [37] employs k-means clustering for query grouping, considering the semantic information of tokens for enhanced feature learning. Nonetheless, the iterative nature of k-means clustering and the potential for uneven token counts per cluster can impact the efficiency of parallel attention operations.
Figure 1: Comparison among Window Partition, Dynamic Group by k-means, and Semantic Equitable Clustering. Our Semantic Equitable Clustering incorporates image semantics while maintaining efficient clustering, eliminating the need for iterative processes such as in k-means. Furthermore, it enables equi-partitioning of tokens, promoting efficient GPU processing without necessitating additional CUDA optimization.
Given these considerations, an optimal token partitioning scheme should efficiently segregate tokens, incorporate semantic information, and efficiently utilize computational resources (e.g., GPU). In response, we introduce a simple, fast and equitable clustering approach named Semantic Equitable Clustering (SEC). SEC segments tokens based on their relevance to global semantic information. Specifically, we employ global pooling to generate a global token encapsulating global semantic information. The similarity between this global token and all other tokens is then computed, reflecting global semantic relevance. Upon obtaining the similarity matrix, tokens (excluding the global token) are sorted by similarity scores, and the tokens with similar scores are grouped into clusters, ensuring uniform token distribution across clusters. As depicted in Fig. 1, SEC comprehensively considers token semantics and completes the clustering process in a single iteration, unlike the multi-iteration k-means. The resulting clusters, containing an equal number of tokens, can be processed in parallel by the GPU efficiently.
Figure 2: Left: Top-1 accuracy v.s. FLOPs on ImageNet-1K of recent SOTA models. Right: Comparison among different vision language connectors on LLaVA-1.5
Building upon Semantic Equitable Clustering (SEC), we introduce the Semantic Equitable Clustering Vision Transformer (SECViT), a versatile vision backbone that is adaptable to a wide spectrum of downstream tasks. As demonstrated in Fig. 2, SECViT exhibits significant performance improvements compared to previous state-of-the-art (SOTA) models. Impressively, SECViT attains an accuracy of 84.3% utilizing merely 4.6GFLOPS, without the need for additional training data or supervision. This superior performance is maintained across different model scales. Furthermore, SECViT proves its proficiency in downstream tasks, including but not limited to, object detection, instance segmentation, and semantic segmentation.
Beyond vision tasks, we also apply SEC to multimodal large language models (MLLM) such as LLaVA-1.5 [36] to serve as an efficient vision language connector. Specifically, we use SEC to cluster the vision tokens, and then merge all the tokens at corresponding positions within each cluster into a single token. Experiments demonstrate that this approach significantly enhances the efficiency of LLaVA-1.5 while improving the model’s performance.
2 Related Works
Vision Transformer.
The Vision Transformer (ViT) [11] is considered a powerful visual architecture. Many works have improved the Vision Transformer, including enhancing its training efficiency and reducing its computational cost [38, 10, 43, 24, 61]. DeiT [43] uses distillation loss and incorporates extensive data augmentation methods into the ViT training process. Hierarchical structures represented by PVT [45, 46, 41, 16, 51] reduce the number of tokens in global attention by downsampling the keys and values (KV), thereby low the computational cost. In addition to them, some methods directly prune tokens based on their importance, retaining important tokens [40, 32]. This reduces the number of tokens and subsequently lowers the computational cost of the model. Another highly representative approach is to group all tokens such that each token can only attend to tokens within its own group [38, 10, 61, 9, 37]. This method also significantly reduces the computational cost of self-attention.
Grouping-Based Vision Transformer.
Most grouping-based attention mechanisms perform grouping based on spatial structure [38, 10, 44, 9, 37]. Specifically, the Swin-Transformer [38] divides all tokens into equally sized windows based on their spatial positions, where each token can only attend to tokens within its own window. This significantly reduces the model’s computational cost. In addition to dividing tokens into small windows along the spatial dimension, DaViT [9] also splits channels into multiple groups along the channel dimension. Unlike the above methods that only consider positional information for grouping, DGT [37] takes semantic information into account by using k-means clustering to group the queries.
Vision Language Connector.
The vision language connector is a critical component in MLLMs [36, 3, 23]. It aligns vision tokens with language tokens. Typical vision language connectors include MLP [36], Resampler [1], C-Abstractor [3], and others. Although MLP performs well, it introduces a significant number of vision tokens, which hampers the model’s efficiency. On the other hand, connectors like Resampler improve the model’s efficiency, but at the cost of reduced performance. Unlike these methods, our proposed SEC consider the semantic information of each token and significantly enhances the model’s efficiency while maintaining its performance.
Figure 3: (a) Illustration of SECViT (b) Applying SEC to vision language connector. (c) Illustration of Semantic Equitable Clustering for ViT and vision language connector.
3 Method
3.1 Overall Architecture
The overall architecture of SECViT is shown in Fig. 3(a). SECViT consists of four stages with downsampling factors of $\frac{1}{4}$, $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$, respectively. This structural design facilitates downstream tasks, such as object detection, in constructing feature pyramids. A SECViT block is composed of three modules. For each block, the input tensor $X_{in}\in\mathbb{R}^{C\times H\times W}$ is fed into the CPE to introduce the positional information. Then, The Self-Attention based on the Semantic Equitable Clustering (SEC) is employed to serve as the token mixer. The final FFN is utilized to integrate channel-wise information of tokens.
Beyond the design of the backbone, we also utilize SEC in the design of the vision language connector in MLLM [36]. For the vision tokens output by ViT, we use SEC to cluster the tokens. For each position corresponding to a cluster, we use attentive pooling to merge them into a single token, thereby reducing the number of vision tokens. The process is shown in Fig. 3(b).
3.2 Semantic Equitable Clustering
As previously mentioned, the design objectives of Semantic Equitable Clustering are threefold: 1) Fully consider the semantic information contained in different tokens during clustering. 2) Unlike k-means and other clustering methods that require multiple iterations, Semantic Equitable Clustering can complete clustering in a single step. 3) Ensure an equal number of tokens in each cluster to facilitate parallel processing on GPUs. In the following paragraphs, we will describe in detail how our Semantic Equitable Clustering achieves these three objectives. And the whole process is illustrated in the Fig. 3(c).
Single Clustering Center Related to Semantics.
k-means is relatively complex for two reasons. First, it has multiple cluster centers, and each token needs to calculate its distance to each cluster center to determine its cluster membership. Second, the determination of each cluster center in k-means is not precise and requires multiple iterations to accurately establish the cluster centers.
To address these two issues, we first discard the use of multiple cluster centers and instead calculate the distance between each token and a single center. Based on each token’s distance to this center, we divide the tokens into different intervals. Then, to ensure that our chosen center contains the most comprehensive semantic information, we directly use the result of average pooling of all tokens as the center token. This is because, in most vision foundation models, the output of the average pool is assumed to contain the richest semantic information and is thus used for classification [38, 10, 6, 12]. Specifically, the process for determining the cluster center is shown in Eq. 1:
$\centering\begin{split}Q=W_{Q}X,&K=W_{K}X,V=W_{V}X,\\ &k_{c}={\rm Pool}(K).\end{split}\@add@centering$ (1)
Where $W_{K}$ is a learnable matrix. $k_{c}$ is the determined cluster center. $X$ is the set of input tokens.
Distance Metric Suitable for ViT.
Unlike the Euclidean distance calculation used in the k-means algorithm for computing the distance between tokens, during the actual computation of Self-Attention, similarity between query and key is computed through dot product. To better adapt to the characteristics of Self-Attention, we also measure the distance between tokens using a method similar to dot product. Specifically, we calculate the cosine similarity between the cluster center and each token, and then sort the tokens according to the magnitude of the computed results. The specific process is shown in Eq. 2:
$\centering\begin{split}&sim=\frac{K\cdot k_{c}}{||K||\cdot||k_{c}||},\\ &idx={\rm argsort}(sim),\\ Q^{*}=Q[idx]&,K^{*}=K[idx],V^{*}=V[idx].\end{split}\@add@centering$ (2)
Where $sim$ is the similarity matrix between $K$ and $k_{c}$, the ${\rm argsort}(sim)$ returns the indices of $sim$ sorted in descending order. $Q^{*},K^{*},V^{*}$ are $Q,K,V$ rearranged according to ${\rm argsort}(sim)$.
Equally Partition Tokens based on Distance.
The obtained $Q^{*}$, $K^{*}$, and $V^{*}$ from the previous step have been sorted based on their distances to the cluster center. For the design of vision backbone, we directly group them, so tokens with similar distances to the cluster center are classified into the same cluster. This allows us to directly control an equal number of tokens in each cluster. This process can be clearly illustrated in Fig. 3(c) and denoted as follows:
$\centering\begin{split}Q_{m}=Q^{*}[{m\times N:\left(m+1\right)N}],\\ K_{m}=K^{*}[{m\times N:\left(m+1\right)N}],\\ V_{m}=V^{*}[{m\times N:\left(m+1\right)N}].\end{split}\@add@centering$ (3)
where $N$ is the basic token number of each cluster for equal partition and $m$ is the index of the cluster
Based on the above steps, we have completed the clustering process that captures semantic information in the image with minimal sorting cost. Moreover, compared to k-means, we have achieved equi-partitioning of each cluster. After clustering is completed, we apply standard Self-Attention to the tokens within each cluster, thereby completing the interaction of information between tokens:
$\centering Y_{m}=\text{Attn}\left(Q_{m},K_{m},V_{m}\right).\@add@centering$ (4)
For the design of vision language connector, we group the tokens according to their similarity, and the tokens within each group are interleaved, as shown in Eq. 5:
$\centering\begin{split}Q_{n}=Q^{*}[n:N:L],\\ K_{n}=K^{*}[n:N:L],\\ V_{n}=V^{*}[n:N:L].\end{split}\@add@centering$ (5)
in which $L$ is the token’s sequence length, $n$ is the index of group tokens. $N$ is the basic token number of each cluster. After obtaining the token groups, we perform pooling on $Q$ to effectively reduce the number of tokens input to the LLM, with each group’s output becoming a single token, as shown in Eq 6.
$Y_{n}={\rm Attn}({\rm Pool}(Q_{n}),K_{n},V_{n}).$ (6)
3.3 Difference between SEC and EViT.
We use the most representative example, EViT [32], to illustrate the differences between SEC and other methods based on the similarity between the global token and other tokens.
Pruning v.s. Clustering. Most similarity-based methods, such as EViT, are pruning methods, where tokens with low similarity to the [cls] token are merged during the forward process, thereby reducing the number of tokens and decreasing computational cost. In contrast, our proposed SECViT employs a clustering-based approach, performing attention operations within each cluster.
The role of the [cls] token. In methods like EViT, the [cls] token serves as a measure of importance of a token. Each token computes its similarity to the [cls] token, with higher similarity tokens deemed more important. The less important tokens are abandoned. In contrast, in SEC, the [cls] token (obtained by average pooling over all tokens) measures similarity between tokens. Each token computes its similarity score to the [cls] token; tokens with similar scores are considered to be more similar and grouped into one cluster. Attention is calculated only within the same cluster.
4 Experiments
We first make strict comparison with hierarchical/plain baselines. Then we conduct experiments on a wide range of vision tasks for SECViT, including image classification, object detection, instance segmentation, and semantic segmentation. We also verify the role of SEC in MLLM based on LLaVA-1.5 [36]. More details, experiments, and comparison of models’ efficiency can be found in the Appendix.
4.1 SEC for vision models
Model Params (M) FLOPs (G) Throughput (imgs/s) Acc $AP^{b}$ $AP^{m}$ mIoU
DeiT-S [43] 22 4.6 3204 79.8 44.5 40.1 43.0
EViT-DeiT-S (keeprate=0.9) 22 4.0 3428 79.8 not suit not suit not suit
SEC-DeiT-S (num_cluster=4) 22 4.1 3412 80.5 (+0.7) 47.7 (+3.2) 42.7 (+2.6) 47.5 (+4.5)
DeiT-B 86 17.6 1502 81.8
SEC-DeiT-B 86 14.8 1682 82.4 (+0.6)
Swin-T 29 4.5 1723 81.3 43.7 39.8 44.5
SEC-Swin-T 29 4.8 1482 83.8 (+2.5) 48.3 (+4.6) 43.4 (+3.6) 49.3 (+4.8)
Swin-S 50 8.8 1006 83.0 45.7 41.1 47.6
SEC-Swin-S 50 9.2 804 85.0 (+2.0) 50.2 (+4.5) 44.7 (+3.6) 51.3 (+3.7)
Table 1: Comparison with Hierarchy/Plain baselines. Inference speed are measured on the A100 GPU.
Model Params (M) FLOPs (G) Method Pretrain epoch Acc(%)
Swin-B [38] 88 15.4 Supervised 83.5
ConvNeXt V2-B [50] 88 15.4 Supervised 84.3
SEC-Swin-B 88 16.2 Supervised 85.3
Swin-B [38] 88 15.4 SimMIM [53] 800 84.0(+0.5)
ConvNeXt V2-B [50] 88 15.4 FCMAE [50] 800 84.6(+0.3)
SEC-Swin-B 88 16.2 SimMIM [53] 800 85.9(+0.6)
Table 2: Comparison with baselines on self-supervised setting.
Cost Model Parmas (M) FLOPs (G) Top1-acc (%)
tiny model $\sim 2.5$G PVTv2-b1 [46] 13 2.1 78.7
TCFormer-light [59] 14 3.8 79.4
QuadTree-B-b1 [42] 14 2.3 80.0
MPViT-XS [28] 11 2.9 80.9
BiFormer-T [61] 13 2.2 81.4
CrossFormer-T [47] 28 2.9 81.5
FAT-B2 [12] 14 2.0 81.9
GC-ViT-XT [20] 20 2.6 82.0
SMT-T [34] 12 2.4 82.2
RMT-T [13] 14 2.5 82.4
SECViT-T 15 2.5 82.7
small model $\sim 4.5$G PS-ViT-B14 [58] 21 5.4 81.7
DVT-T2T-ViT-19 [49] 39 6.2 81.9
ConvNeXt-T [39] 29 4.5 82.1
TCFormer [59] 26 5.8 82.3
SG-Former-S [15] 23 4.8 83.2
StructViT-S-8-1 [25] 24 5.4 83.3
InternImage-T [48] 30 5.0 83.5
MLLA-T [18] 25 4.2 83.5
MaxViT-T [44] 31 5.6 83.6
FAT-B3 [12] 29 4.4 83.6
SMT-S [34] 20 4.8 83.7
BiFormer-S [61] 26 4.5 83.8
SECViT-S 27 4.6 84.3
base model $\sim 9.0$G ConvNeXt-S [39] 50 8.7 83.1
NAT-S [19] 51 7.8 83.7
Quadtree-B-b4 [42] 64 11.5 84.0
MOAT-1 [54] 42 9.1 84.2
InternImage-S [48] 50 8.0 84.2
GC-ViT-S [20] 51 8.5 84.3
BiFormer-B [61] 57 9.8 84.3
iFormer-B [41] 48 9.4 84.6
FAT-B4 [12] 52 9.3 84.8
SECViT-B 57 9.8 85.2
large model $\sim 18.0$G CrossFormer-L [47] 92 16.1 84.0
SMT-L [34] 81 17.7 84.6
DaViT-B [9] 88 15.5 84.6
SG-Former-B [15] 78 15.6 84.7
iFormer-L [41] 87 14.0 84.8
InterImage-B [48] 97 16.0 84.9
GC-ViT-B [20] 90 14.8 85.0
RMT-L [13] 95 18.2 85.5
SECViT-L 101 18.2 85.7
Table 3: Comparison with the state-of-the-art on ImageNet-1K classification.
Backbone Params (M) FLOPs (G) Mask R-CNN 1×1$\times$1 × Params (M) FLOPs (G) RetinaNet 1×1$\times$1 ×
$AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$ $AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{b}_{S}$ $AP^{b}_{M}$ $AP^{b}_{L}$
PVTv2-B1 [46] 33 243 41.8 54.3 45.9 38.8 61.2 41.6 23 225 41.2 61.9 43.9 25.4 44.5 54.3
FAT-B2 [12] 33 215 45.2 67.9 49.0 41.3 64.6 44.0 23 196 44.0 65.2 47.2 27.5 47.7 58.8
RMT-T [13] 33 218 47.1 68.8 51.7 42.6 65.8 45.9 23 199 45.1 66.2 48.1 28.8 48.9 61.1
SECViT-T 34 221 47.8 69.5 52.5 43.0 66.7 46.3 24 202 45.8 66.8 49.2 29.1 49.8 60.9
MPViT-S [28] 43 268 46.4 68.6 51.2 42.4 65.6 45.7 32 248 45.7 57.3 48.8 28.7 49.7 59.2
MLLA-T [18] 44 255 46.8 69.5 51.5 42.1 66.4 45.0
STViT-S [22] 44 252 47.6 70.0 52.3 43.1 66.8 46.5
RMT-S [13] 46 262 49.0 70.8 53.9 43.9 67.8 47.4 36 244 47.8 69.1 51.8 32.1 51.8 63.5
SECViT-S 45 262 49.9 70.9 54.7 44.6 68.3 47.7 35 240 48.4 69.4 52.0 31.3 53.3 63.8
ScalableViT-B [56] 95 349 46.8 68.7 51.5 42.5 65.8 45.9 85 330 45.8 67.3 49.2 29.9 49.5 61.0
InternImage-S [48] 69 340 47.8 69.8 52.8 43.3 67.1 46.7
MLLA-S [18] 63 319 49.2 71.5 53.9 44.2 68.5 47.2
STViT-B [22] 70 359 49.7 71.7 54.7 44.8 68.9 48.7
SECViT-B 76 371 51.5 72.9 56.7 45.4 69.9 48.7 63 349 49.3 70.3 52.9 32.0 53.8 64.8
Focal-B [55] 110 533 47.8 70.2 52.5 43.2 67.3 46.5 101 514 46.3 68.0 49.8 31.7 50.4 60.8
CSwin-B [10] 97 526 48.7 70.4 53.9 43.9 67.8 47.3
InternImage-B [48] 115 501 48.8 70.9 54.0 44.0 67.8 47.4
MLLA-B [18] 115 502 50.5 72.0 55.4 45.0 69.3 48.6
SECViT-L 119 550 52.0 73.5 57.3 46.3 70.6 49.8 105 527 50.2 71.4 53.9 33.2 54.5 66.3
Table 4: Comparison to other backbones using ”1×1$\times$1 ד schedule on COCO.
Backbone Params (M) FLOPs (G) $AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
Mask R-CNN 3×3$\times$3 ×+MS
GC-ViT-T [20] 48 291 47.9 70.1 52.8 43.2 67.0 46.7
MLLA-T [18] 44 255 48.8 71.0 53.6 43.8 68.0 46.8
SMT-S [34] 40 265 49.0 70.1 53.4 43.4 67.3 46.7
InternImage-T [48] 49 270 49.1 70.4 54.1 43.7 67.3 47.3
RMT-S [13] 46 262 50.7 71.9 55.6 44.9 69.1 48.4
SECViT-S 45 262 51.6 72.5 55.9 45.6 69.9 48.8
NAT-S [19] 70 330 48.4 69.8 53.2 43.2 66.9 46.4
InternImage-S [48] 69 340 49.7 71.1 54.5 44.5 68.5 47.8
SMT-B [34] 52 328 49.8 71.0 54.4 44.0 68.0 47.3
MLLA-S [18] 63 319 50.5 71.8 55.2 44.9 69.1 48.2
RMT-B [13] 73 373 52.2 72.9 57.0 46.1 70.4 49.9
SECViT-B 75 371 52.8 73.6 57.7 46.4 70.8 49.9
Cascade Mask R-CNN 3×3$\times$3 ×+MS
GC-ViT-T [20] 85 770 51.6 70.4 56.1 44.6 67.8 48.3
SMT-S [34] 78 744 51.9 70.5 56.3 44.7 67.8 48.6
UniFormer-S [30] 79 747 52.1 71.1 56.6 45.2 68.3 48.9
RMT-S [13] 83 741 53.2 72.0 57.8 46.1 69.8 49.8
SECViT-S 83 741 54.1 72.8 58.6 47.0 70.3 51.0
NAT-S [19] 108 809 51.9 70.4 56.2 44.9 68.2 48.6
GC-ViT-S [20] 108 866 52.4 71.0 57.1 45.4 68.5 49.3
CSWin-S [10] 92 820 53.7 72.2 58.4 46.4 69.6 50.6
UniFormer-B [30] 107 878 53.8 72.8 58.5 46.4 69.9 50.4
RMT-B [13] 111 852 54.5 72.8 59.0 47.2 70.5 51.4
SECViT-B 114 849 55.4 74.1 59.9 47.8 71.7 51.7
Table 5: Comparison with other backbones using ”$3\times+\mathrm{MS}$“ schedule on COCO.
Model Semantic FPN 80K Upernet 160K
Params (M) FLOPs (G) mIoU (%) Params (M) FLOPs (G) mIoUss (%)
PVTv2-B1 [46] 18 136 42.5
VAN-B1 [17] 18 140 42.9
RMT-T [13] 17 136 46.4
SECViT-T 18 136 47.2 44 894 48.8
StructViT-S [25] 26 271 46.9
MogaNet-S [31] 29 189 47.7 55 946 49.2
SMT-S [34] 50 935 49.2
SGFormer-S [15] 25 205 49.0 52.5 989 49.9
RMT-S [13] 30 180 49.4 56 937 49.8
SECViT-S 30 180 49.6 56 936 50.6
MogaNet-B [31] 74 1050 50.1
InterImage-S [48] 80 1017 50.2
StructViT-B [25] 54 529 48.5
RMT-B [13] 57 294 50.4 83 1051 52.0
SECViT-B 60 291 50.7 86 1048 52.2
MogaNet-L [31] 113 1176 50.9
MLLA-B [18] 128 1183 51.9
SGFormer-B [15] 81 475 50.6 109 1304 52.0
RMT-L [13] 98 482 51.4 125 1241 52.8
SECViT-L 103 475 52.2 131 1256 53.8
Table 6: Comparison with the state-of-the-art on ADE20K.
Model Connector V-T Num Time Speed TextVQA GQA VQAv2 POPE MME
LLaVA-1.5 MLP 576+1 194s 1.0×1.0$\times$1.0 × 58.2 62.0 78.5 86.1 1510.7
LLaVA-1.5+Resampler Resampler 288+1 126s 1.5×1.5$\times$1.5 × 52.1 56.8 76.0 83.1 1393.2
LLaVA-1.5+EViT MLP+EViT 288+1 126s 1.5×1.5$\times$1.5 × 54.6 60.0 77.9 84.3 1483.2
LLaVA-1.5+SEC MLP+SEC 288+1 126s 1.5×1.5$\times$1.5 × 60.1 63.5 78.9 87.7 1510.7
LLaVA-1.5+Resampler Resampler 256+1 116s 1.7×1.7$\times$1.7 × 51.6 56.0 75.2 82.7 1387.2
LLaVA-1.5+Pool MLP+Pool 256+1 116s 1.7×1.7$\times$1.7 × 52.4 57.6 76.4 83.3 1415.5
LLaVA-1.5+EViT MLP+EViT 256+1 116s 1.7×1.7$\times$1.7 × 52.8 59.6 77.1 83.7 1443.7
LLaVA-1.5+SEC MLP+SEC 256+1 116s 1.7×1.7$\times$1.7 × 59.6 63.2 78.6 87.1 1505.2
LLaVA-1.5+Resampler Resampler 192+1 102s 1.9×1.9$\times$1.9 × 50.1 55.2 74.3 82.7 1337.6
LLaVA-1.5+EViT MLP+EViT 192+1 102s 1.9×1.9$\times$1.9 × 51.6 58.6 76.3 83.1 1427.6
LLaVA-1.5+SEC MLP+SEC 192+1 102s 1.9×1.9$\times$1.9 × 57.7 62.7 78.4 86.7 1500.1
LLaVA-1.5+Resampler Resampler 144+1 94s 2.1×2.1$\times$2.1 × 47.6 54.6 72.0 81.9 1293.7
LLaVA-1.5+Pool MLP+Pool 144+1 94s 2.1×2.1$\times$2.1 × 50.0 56.2 73.6 81.9 1310.7
LLaVA-1.5+EViT MLP+EViT 144+1 94s 2.1×2.1$\times$2.1 × 51.2 58.0 76.0 83.1 1393.6
LLaVA-1.5+SEC MLP+SEC 144+1 94s 2.1×2.1$\times$2.1 × 56.8 62.0 78.0 86.1 1487.1
Table 7: Comparison of different vision language connectors on LLaVA-1.5. “V-T Num” denotes the quantity of visual tokens. The computation expense is impacted by V-T Num, with larger values resulting in higher costs. “Speed” refers to the comparative inference velocity relative to LLaVA-1.5. “Time” is the average inference time. Inference speed are measured on the A100.
Model LLM Connector V-T Num Res TextVQA GQA VQAv2 VisWiz SQAimg Speed ($\uparrow$)
7B LLM
Shikra [5] Vicuna-7B MLP 257 224 - - 77.4 - - -
IDEFICS-9B [27] LLaMA-7B Cross Attn 257 224 - 38.4 50.9 35.5 - -
Qwen-VL [1] Qwen-7B Resampler 256 448 - 59.3 78.8 35.2 67.1 -
Qwen-VL-Chat [1] Qwen-7B Resampler 256 448 - 57.5 78.2 38.9 68.2 -
LLaVA-1.5 [35] Vicuna-7B MLP 577 336 58.2 62.0 78.5 50.0 66.8 1.0×1.0$\times$1.0 ×
LLaVA-1.5+SEC (ours) Vicuna-7B MLP+SEC 257 336 59.6 63.2 78.9 52.8 69.6 $\mathbf{1.7\times}$
13B LLM
InstructBLIP [8] Vicuna-13B Q-Former 32 224 - 49.5 - 33.4 63.1 -
BLIP-2 [29] Vicuna-13B Q-Former 32 224 - 41.0 41.0 19.5 61.0 -
LLaVA-1.5 [35] Vicuna-13B MLP 577 336 61.2 63.3 80.0 53.6 71.6 1.0×1.0$\times$1.0 ×
LLaVA1.5+SEC (ours) Vicuna-13B MLP+SEC 257 336 62.3 64.3 80.0 54.7 72.0 $\mathbf{1.8\times}$
Table 8: Results on General VQA tasks.
Model LLM Connector V-T Num Res POPE MMB MM-Vet Speed ($\uparrow$)
7B LLM
MiniGPT-4 [60] Vicuna-7B Resampler 32 224 72.2 24.3 22.1 -
mPLUG-Owl2 [57] LLaMA2-7B Resampler 32 224 - 49.4 - -
LLaMA-AdapterV2 [14] LLaMA2-7B LLaMA-Adapter 257 224 - 41.0 31.4 -
Shikra [5] Vicuna-7B MLP 257 224 - 58.8 - -
Qwen-VL [1] Qwen-7B Resampler 256 448 - 38.2 - -
Qwen-VL-Chat [1] Qwen-7B Resampler 256 448 - 60.6 - -
LLaVA-1.5 [35] Vicuna-7B MLP 577 336 86.1 64.3 31.1 1.0×1.0$\times$1.0 ×
LLaVA1.5+SEC (ours) Vicuna-7B MLP+SEC 145 336 86.1 68.4 31.7 $\mathbf{2.1\times}$
13B LLM
MiniGPT-4 [60] Vicuna-13B Resampler 32 224 - - 24.4 -
BLIP-2 [29] Vicuna-13B Q-Former 32 224 85.3 - 22.4 -
LLaVA-1.5 [35] Vicuna-13B MLP 577 336 86.2 67.7 36.1 1.0×1.0$\times$1.0 ×
LLaVA-1.5+SEC (ours) Vicuna-13B MLP+SEC 145 336 86.4 69.2 37.3 $\mathbf{2.2\times}$
Table 9: Results on benchmark designed for MLLMs.
Strict Comparison with Baselines.
We select two baselines: hierarchical backbone Swin-Transformer [38] and plain backbone DeiT [43] to make a comparison with our SEC based model. In the comparison models (SEC-Swin and SEC-DeiT), we merely substitute the attention mechanism in the original model with our SEC based Self-Attention and without introducing any other modules. As shown in Tab. 1, we conduct experiments on image classification, object detection, insatance segmentation and semantic segmentation, the simple replacement of the attention mechanism yields significant advantages in both performance and efficiency.
In addition to the supervised scenario, we also train the model with SimMIM [53] in the self-supervised scenario. As shown in Tab. 2, SEC also performs exceptionally well in the self-supervised scenario.
Image Classification.
We compare our SECViT with numerous state-of-the-art models, the results are shown in Tab.3. We adopt the training strategy proposed in DeiT [43], with the only supervision is cross entropy loss. All of our models are trained from scratch for 300 epochs with the input resolution of $224\times 224$. SECViT consistently outperforms preceding models across all scales. Notably, SECViT-S attains a Top1-accuracy of 84.3% with a mere 27M parameters and 4.6G FLOPs. The comparison of the models’ efficiency can be found in Appendix.
Object Detection and Instance Segmentation.
We utilize MMDetection [4] to implement Mask-RCNN [21], Cascade Mask R-CNN [2], and RetinaNet [33] to evaluate the performance of the SECViT. Tab. 5 and Tab. 4 show the results of SECViT with different detection frameworks. The results show that SECViT performs better than its counterparts in all comparisons.
Semantic Segmentation.
We utilize Semantic FPN [26] and UperNet [52] to validate our SECViT’s performance, implementing these frameworks via MMSegmentation [7]. The results of semantic segmentation can be found in the Tab. 6. All the FLOPs are measured with the input resolution of $512\times 2048$. SECViT achieves the best performance in all settings.
4.2 SEC for MLLM
SEC can greatly facilitate the design of vision language connectors in MLLMs. First, we conduct a rigorous comparison between SEC and various baseline vision language connectors based on LLaVA-1.5. Then, we compare LLaVA-1.5+SEC with several popular contemporary MLLMs.
Strict Comparison with Baselines.
In Tab. 7, we strictly compare various commonly used vision language connectors, including MLP, Resampler [1], Pooling, and EViT [32], which has achieved success in the design of ViT. Among these, MLP is the original design in LLaVA-1.5 [36], capable of achieving good results. However, it incurs significant computational cost due to the excessive vision tokens. To address this issue, some connectors attempt to use fewer vision tokens to accelerate LLaVA-1.5. Nonetheless, these adjustments inevitably lead to performance degradation. The results in Tab. 7 show that using SEC can effectively accelerate the inference of LLaVA-1.5 without causing performance degradation, and can even improve the performance of LLaVA-1.5 to a certain extent.
Comparison with Popular MLLMs.
In Tab. 8 and Tab. 9, we compare LLaVA-1.5 equipped with SEC as a vision-language connector with other MLLMs. It is evident that SEC not only enhances the performance of MLLMs across various benchmarks but also significantly improves the efficiency of the models. This fully demonstrates the effectiveness of SEC in extracting visual information.
4.3 Ablation Study
In this section, we present some of the ablation study results for SEC, and more results can be found in the Appendix.
Number of Vision Tokens in Each Clusters.
The number of vision tokens has a significant impact on the performance and speed of the model. We thoroughly investigate the effect of the number of vision tokens on SECViT. As shown in Tab. 10, the number of vision tokens in each cluster greatly influences the model’s performance. Specifically, in downstream dense prediction tasks, having too few tokens in each cluster leads to substantial performance degradation. When the number of tokens in each cluster is too large, the model’s performance does not see a significant improvement, but its speed decreases.
V-T num Params (M) FLOPs (G) Throughput (imgs/s) Acc $AP^{b}$ $AP^{m}$ mIoU
98 15 2.5 2004 82.7 47.8 43.0 47.2
196 15 3.1 1722 83.0 (+0.3) 48.2 (+0.4) 43.4 (+0.4) 47.5 (+0.3)
64 15 2.5 1946 82.7 (+0.0) 47.8 (+0.0) 42.8 (-0.2) 46.9 (-0.3)
49 15 2.4 2102 82.6 (-0.1) 47.5 (-0.3) 42.7 (-0.3) 47.7 (-0.5)
24 15 2.3 2186 82.0 (-0.7) 45.9 (-1.9) 40.6 (-2.4) 44.6 (-2.6)
Table 10: Effect of the number of vision tokens in each cluster. “V-T num” means the number of vision tokens in each cluster. The experiments are conducted based on SECViT-T.
4.4 Visualization of SEC
To further understand the working mechanism of SEC, we visualize some clustering results for SECViT. As shown in Fig. 5, the left side presents the clustering results of vision tokens at different stages of the model. From the clustering results, we analyze that in the shallow layers, the model distinguishes fine-grained features well, while in the deeper layers, it captures global semantic features effectively. The right side shows the Grad-CAM diagrams at different stages of the model, from which we can draw similar conclusions to the clustering results. More visualization results can be found in Appendix.
Figure 5: Visualization for SEC.
Number of Vision Tokens Outputs by SEC.
MLLM is quite sensitive to the number of vision tokens. We conduct a detailed exploration based on LLaVA-1.5 regarding the number of vision tokens output by SEC, as shown in Tab. 11. The first row represents the speed and performance of the original LLaVA-1.5 without using SEC. Compared to LLaVA-1.5, employing SEC effectively reduces the number of vision tokens and improves training efficiency. As the number of vision tokens decreases, the model’s performance shows a slight decline, but its efficiency is further enhanced.
V-T num Time Speed TextVQA GQA VQAv2 POPE MM-Vet
576+1 21h 1.0×1.0$\times$1.0 × 58.2 62.0 78.5 86.1 31.1
288+1 14h 1.5×1.5$\times$1.5 × 60.1(+1.9) 63.5(+1.5) 78.9(+0.4) 87.7(+1.6) 33.2(+2.1)
256+1 13h 1.6×1.6$\times$1.6 × 59.6(+1.4) 63.2(+0.3) 78.6(+0.1) 87.1(+1.0) 32.7(+1.6)
192+1 11h 1.9×1.9$\times$1.9 × 57.7(-0.5) 62.7(+0.7) 78.4(-0.1) 86.7(+0.6) 32.1(+1.0)
144+1 10h 2.1×2.1$\times$2.1 × 56.8(-1.4) 62.0(+0.0) 78.0(-0.5) 86.1(+0.0) 31.7(+0.6)
Table 11: Effect of the number of vision tokens outputs by SEC. “V-T num” means the number of vision tokens output by SEC. The experiments are conducted based on LLaVA-1.5 [36].
5 Conclusion
We propose a simple and straightforward clustering method for vision tokens—Semantic Equitable Clustering (SEC). This method assigns each token to a cluster by calculating the similarity between each token and a global token, and completes the whole clustering process in only one step. Our clustering method takes into account the semantic information contained in the tokens, and ensures an equal number of tokens in each cluster, facilitating efficient parallel processing on modern GPUs. Based on Semantic Equitable Clustering, we designed SECViT, a versatile vision backbone that achieves impressive results across various vision tasks, including image classification, object detection, instance segmentation, and semantic segmentation. Besides, SEC can also be conveniently applied to multimodal large language models (MLLM) to serve as a vision language connector and benefits the model’s efficiency.
6 Acknowledgements
This work is partially funded by Beijing Natural Science Foundation (4252054), Youth Innovation Promotion Association CAS(Grant No.2022132), Beijing Nova Program(20230484276), and CCF-Kuaishou Large Model Explorer Fund (NO. CCF-KuaiShou 2024005).
[…]
Supplementary Material
Appendix A Experimental Details
SECViT’s Architectures.
SECViT’s architecture details are illustrated in Table 12. In SECViT, we adopt four $3\times 3$ convolutions to embed the input image into tokens, batch normalization and GELU are used after each convolution. $3\times 3$ convolutions with stride 2 are used between stages to reduce the feature resolution. $3\times 3$ DWConvs are adopted in CPE. For all models, we set the number of clusters in the first three stages to 32, 8, and 2, respectively.
Model Blocks Channels Heads Ratios Params(M) FLOPs(G)
SECViT-T [2, 2, 9, 2] [64, 128, 256, 512] [2, 4, 8, 16] 3 15 2.5
SECViT-S [4, 4, 18, 4] [64, 128, 256, 512] [2, 4, 8, 16] 3 27 4.6
SECViT-B [4, 8, 26, 9] [80, 160, 320, 512] [2, 4, 8, 16] 3 57 9.8
SECViT-L [4, 8, 26, 9] [112, 224, 448, 640] [4, 8, 14, 20] 3 101 18.2
SECViT-XL [6, 12, 28, 12] [128, 256, 512, 1024] [4, 8, 16, 32] 3 205 36.4
Table 12: Detailed Architectures of our models.
Appendix B More experimental Results
Efficiency Comparison.
In Tab. 13, we compare the inference efficiency of various models in detail. From this, we can see that the ViT based on SEC demonstrates the best performance-speed tradeoff.
Model Params(M) FLOPs(G) Throughput(imgs/s) Top1-Acc(%)
DeiT-S [43] 22 4.6 3204 79.8
EViT-DeiT-S (keeprate=0.9) [32] 22 4.0 3428 79.8
SEC-DeiT-S (num_cluster=4) 22 4.1 3412 80.5
DeiT-B [43] 86 17.6 1502 81.8
SEC-DeiT-B 86 14.8 1682 82.4
PVTv2-b1 [46] 13 2.1 2204 78.7
TCFormer-light [59] 14 3.8 417 79.4
MPViT-XS [28] 11 2.9 1496 80.9
BiFormer-T [61] 13 2.2 1634 81.4
CMT-XS [16] 15 1.5 1476 81.8
GC-ViT-XT [20] 20 2.6 1308 82.0
SMT-T [34] 12 2.4 638 82.2
RMT-T [13] 14 2.5 1438 82.4
SECViT-T 15 2.5 2004 82.7
Swin-T [38] 29 4.5 1723 81.3
PS-ViT-B14 [58] 21 5.4 1986 81.7
DVT-T2T-ViT-19 [49] 39 6.2 1268 81.9
SGFormer-S [15] 23 4.8 952 83.2
CMT-S [16] 25 4.0 846 83.5
CSwin-S [10] 35 6.9 972 83.6
SMT-S [34] 20 4.8 356 83.7
BiFormer-S [61] 26 4.5 766 83.8
SEC-Swin-T 29 4.8 1482 83.8
SECViT-S 27 4.6 998 84.3
Swin-S [38] 50 8.8 1006 83.0
SGFormer-M [15] 39 7.5 598 84.1
SMT-B [34] 32 7.7 237 84.3
BiFormer-B [61] 57 9.8 498 84.3
MaxViT-S [44] 69 11.7 546 84.5
CMT-B [16] 46 9.3 447 84.5
iFormer-B [41] 48 9.4 688 84.6
RMT-B [13] 54 9.7 430 85.0
SEC-Swin-S 50 9.2 804 85.0
SECViT-B 57 9.8 504 85.2
Swin-B [38] 88 15.5 768 83.5
CSWin-B [10] 78 15.0 660 84.2
SMT-L [34] 80 17.7 158 84.6
SGFormer-B [15] 78 15.6 388 84.7
iFormer-L [41] 87 14.0 410 84.8
MaxViT-B [44] 120 23.4 306 84.9
SEC-Swin-B 88 16.2 696 85.3
SECViT-L 101 18.2 398 85.7
Table 13: Comparison of models’ efficiency. Throughputs are measured on a single A100 with the batch size of 64.
Different Methods for Merging Vision Tokens.
For MLLM, SEC uses an interleaved merge token approach to reduce the number of vision tokens. Conversely, we also explore a sequential merge token method to achieve a similar reduction. The comparison of these two methods is shown in Tab. 14. The direct sequential merge token approach may result in the loss of critical visual information, significantly degrading the model’s performance.
Method V-T num Time Speed TextVQA GQA VQAv2 POPE MM-Vet
Interleaved 288+1 14h 1.5×1.5$\times$1.5 × 60.1 63.5 78.9 87.7 33.2
Sequential 288+1 14h 1.5×1.5$\times$1.5 × 52.8(-7.3) 57.1(-6.2) 75.7(-3.2) 81.7(-6.0) 27.6(-5.6)
Interleaved 144+1 10h 2.1×2.1$\times$2.1 × 56.8 62.0 78.0 86.1 31.7
Sequential 144+1 10h 2.1×2.1$\times$2.1 × 47.2(-9.6) 53.6(-8.4) 71.7(-6.3) 80.0(-6.1) 22.3(-9.6)
Table 14: Different methods for merging vision tokens.
Appendix C More Clustering Results for Complex Scenes
To further illustrate the mechanism of SEC, we visualize more images in complex scenes and their clustering results, as shown in Fig. 6. Specifically, we visualize the clustering results of the first three stages of SECViT. The results further demonstrate that SEC can better learn fine-grained representations in the shallow layers of the model and semantic representations in the deeper layers.
Figure 6: Visualization results for complex scenes.